OpenStudy (anonymous):

need help in math completely confused !! please help! Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

are the x= the answers you tried that didn't work?

OpenStudy (anonymous):

I think x= -9

OpenStudy (anonymous):

−4|x + 5| = −16 |x + 5| = 4 x + 5 = 4 or x + 5 = -4 x = -1 or x = -9

OpenStudy (anonymous):

Thank you so much! mind helping with one more??

OpenStudy (anonymous):

Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution

OpenStudy (anonymous):

This type of math is really confusing if you don't mind explaining as well?

OpenStudy (anonymous):

First, isolate the abs value on one side |x| − 8 = −5 |x| = 3 Now make the left equal the positive right side or the negative right side x = 3 OR x = -3